OpenStudy (sarahc):

Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (jdoe0001):

hmm anyhow, same procedure solve for "t" onthe 1st equation

jimthompson5910 (jim_thompson5910):

We will take advantage of the Pythagorean trig identity to get... \[\Large x^2 + y^2 = (3\cos(t))^2+(3\sin(t))^2\] \[\Large x^2 + y^2 = 9\cos^2(t)+9\sin^2(t)\] \[\Large x^2 + y^2 = 9(\cos^2(t)+\sin^2(t))\] \[\Large x^2 + y^2 = 9(1)\] \[\Large x^2 + y^2 = 9\] Notice how I used the identity \(\Large \cos^2(t)+\sin^2(t)=1\) (which is true for all real numbers t).

OpenStudy (jdoe0001):

actaully.. is a trig.. ok